Circular No. 6757

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970828 J. Greiner and R. Schwarz, Astrophysical Institute, Potsdam; J. Engl-

hauser, Max-Planck-Institut für Extraterrestrische Physik, Garching; and P. J. Groot and T. J. Galama, Astronomical Institute, Amsterdam, report: “A 61 000-s ROSAT observation centered on GRB 970828 (IAUC 6728) during Sept. 3–5 reveals fifteen x-ray sources within the HRI fieldof-view above the 4σ level. One of these, located at α = 18^{h}08^{m}31^{s}.7, δ = +59^{o}18′50′′ (equinox 2000.0; uncertainty ± 10′′; RX J1808.5+5918), is within the RXTE/IPN error box and coincides with the x-ray source detected with ASCA (IAUC 6732). Using a power-law model with photon index 2 and the galactic column N_{H} = 3.7 × 10^{20} cm^{−2}, the observed countrate of (4 ± 1) × 10^{−4} counts/s corresponds to an unabsorbed ROSAT flux (band 0.1–2.4 keV) of 2.5 × 10^{−14} erg cm^{−2} s^{−1}, consistent with a power-law decay. Though the probability of finding an x-ray source of that intensity within an 8-arcmin^{2} error box is rather high (0.002), the coincidence in position and expected fading flux suggests that we detected the x-ray afterglow of GRB 970828. In R-band images obtained with the 4.2m William Herschel Telescope between 3.5 hr and 8 days after the GRB event, we detect seven objects in the ROSAT error box, with R magnitudes in the range 24.1–25.0. None of these is variable by more than 0.3 mag. The object of mag K′ = 19 reported on IAUC 6756 is located within the 10′′ x-ray error circle and coincides with an optical object for which R = 24.9 ± 0.3 (obtained from point-spread-function fitting). The x-ray image, as well as an error box overlay on R- and K-band images, is posted at

http://www.aip.de:8080/~jcg/grb970828.html).”

XTE J1739–302 D. M. Smith, Space Sciences Laboratory, University of California at

Berkeley, communicates: “The position originally reported for XTE J1739– 302 on IAUC 6748 was incorrect. The correct position is 7′ away, at α = 17^{h}38^{m}53^{s}, δ = −30^{o}15′.6 (equinox 2000.0). An error radius of 3′ at 99percent confidence is still appropriate, excluding the original field.”

COMET C/1997 T1 (UTSUNOMIYA) Visual m_{1} estimates: Oct. 14.13 UT, 10.4 (G. W. Kronk, Troy, IL,

0.33-m reflector); 18.76, 9.9 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector); 20.15, 9.4 (C. E. Spratt, Victoria, BC, 0.20-m reflector).

1997 October 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green